# Trial by FIRE: Probing the dark matter density profile of dwarf galaxies with GraphNPE

### v1

@@ 6 Comparison between CDM and SIDM
Figure 11 shows the normalized distribution of fractional residuals for $V_{\mathrm{max}}$ (top) and $M_{\mathrm{200m}}$ (bottom), with CDM and SIDM simulations represented in purple and green, respectively. The mean and standard deviation of each distribution are indicated in the figure.
In the top panel, the mean residuals suggest that GraphNPE tends to overestimate $V_{\mathrm{max}}$ for CDM and SIDM1, consistent with the trends discussed in Section 5. This is further evidenced by the extended tails of the CDM and SIDM1 distributions. The model performs best on the SIDM10 samples, where the $V_{\mathrm{max}}$ residual distribution is more centered around zero. The residual distributions of $M_{\mathrm{200m}}$ in the bottom panel are more comparable between the simulations. The SIDM1 distribution peaks the highest at zero, but also exhibits some outliers toward positive residuals, contributing to a larger overall spread.
Table 6 presents the 68% and 95% coverage fractions for both $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}$ across the CDM and SIDM samples. The reported means and associated errors are derived from 1000 bootstrap resamples. The model shows comparable performance on the CDM and SIDM10 simulations. In contrast, SIDM1 exhibits lower overall coverage, likely driven by the long tail of high residuals shown in Figure 6.
@@ 7 Discussion
Figure 13: The pericentric distance $r_{\mathrm{peri}}$ versus current distance $r_{\mathrm{curr}}$ for Milky Way dwarfs (blue and orange circles) and our FIRE-2 mock dataset (gray contours). The contours are shown at every 10% intervals. Ultra-faint dwarfs $(M_{V}>-7.7)$ are denoted as squares.
@@
Despite differences between the FIRE-2 mock samples and observations, it is important to note that the majority of the observed data samples remain well within the prior distribution. From Table 2, the Plummer scale radius can range from $r_{\mathrm{\star}}\in[20,20000]\,\mathrm{pc}$, [^12] For a Plummer profile, the projected half-stellar mass radius is equal to the scale radius, i.e., $R_{\star,1/2}=r_{\mathrm{\star}}$. while the velocity dispersion can range from $\sigma_{\mathrm{3-D}}\in[0.1,50]\,\mathrm{km/s}$ (or $\sigma_{\mathrm{los}}\in[0.06,28.9]\,\mathrm{km/s}$, assuming isotropy). The three galaxies with the smaller half-stellar mass radii are Triangulum II, Segue 1, and Willman 1 with values of $20.56^{+2.74}_{-2.61}\,\mathrm{pc}$, $23.90^{+3.74}_{-3.40}\,\mathrm{pc}$, and $27.43^{+5.89}_{-5.35}\,\mathrm{pc}$, respectively (see Pace 2024 and references therein). While these technically fall within the prior range, Triangulum II is very close to the lower boundary. This proximity may artificially truncate the posterior distribution, preventing a proper exploration of smaller values and potentially biasing parameter inferences. To better accommodate such observations, we plan to extend the prior limits in future analysis.
@@
Figure 13 shows the pericentric distance $r_{\mathrm{peri}}$ versus the current distance $r_{\mathrm{curr}}$ of the Milky Way dwarfs (circles) and the FIRE-2 simulations (gray contours, shown at every 10% interval. Blue and orange circles represent dwarfs within and outside of the 90% contour, respectively. Similarly, ultra-faint dwarfs $(M_{\mathrm{V}}>-7.7)$ are marked as squares and not expected to be resolved by FIRE-2. The vertical and horizontal dashed lines correspond to $0.2$ and $0.5$ times $R_{\mathrm{200m}}$ of the Milky Way, assumed to be $365\pm 76\,\mathrm{kpc}$ (Deason et al., 2020). Orbital properties are taken from Pace et al. (2022), which uses mixture models to determine the systemic proper motions of the Milky Way satellites and subsequently fits their orbits, accounting for both the Milky Way and LMC potentials. About half of the Milky Way satellites are within the 90% contour of the simulations, which typically occupy the high $r_{\mathrm{peri}}$ and $r_{\mathrm{curr}}$ regions. This difference highlights a numerical limitation: simulated satellites too close to their hosts are often disrupted and become much more challenging to track with rockstar (Behroozi et al., 2013; Mansfield et al., 2024).
@@ 7.2 Comparison with distribution function fitting
First, as with other NPE methods, GraphNPE is amortized, i.e. once trained, the model can be applied to any observation without re-training. In contrast, likelihood-based methods, including DF methods, must be rerun for each new dataset. This significantly limits the parameter space that can be explored with these models, particularly when extending to non-spherical cases. For systems with 1000 tracers, the Agama DF method in Read et al. (2021) takes approximately $\mathcal{O}(10^{3})$ minutes per run, comparable to the total training time of GraphNPE. However, once trained, GraphNPE performs inference in milliseconds, regardless of the tracer counts, making it far more efficient for large-scale applications. While computational cost may not be a major concern for real observations, given that only a few dwarf galaxies have been observed (Pace, 2024), it becomes critical when testing and exploring the modeling parameter space on mock simulations. For instance, the detailed evaluations of mass modeling and tidal effects in Section 5 would be impractical with DF methods on the full FIRE-2 mock dataset, due to their high computational cost.
@@ 8 Conclusion
Dwarf galaxies and their mass density profiles hold the key to understanding structure formation and the particle nature of DM. Accurate mass modeling of these galaxies is essential for constraining DM properties and distinguishing between different theoretical scenarios, such as CDM and SIDM. As new spectroscopic data from surveys such as MUSE-Faint (Zoutendijk et al., 2020, 2021b, 2021a; Júlio et al., 2023) and DELVE (DECam Local Volume Exploration survey; Drlica-Wagner et al. 2022; Tan et al. 2025) continue to push the limits of kinematic measurements to the ultra-faint regime, it is increasingly important to develop robust modeling techniques capable of extracting meaningful constraints even with low stellar tracer counts. Traditional Jeans-based methods often struggle in this regime due to sparse kinematic data, making it critical to adopt approaches that maximize the information extracted from limited observations.
@@
  • When we compare GraphNPE to the GravSphere higher-order Jeans model, we find that the performance is overall similar, with discrepancies primarily arising from differences in modeling choices and prior distributions. Notably, GravSphere tends to favor steeper inner and outer density profiles due to its prior distribution.
@@
Lastly, we briefly compare the performance of GraphNPE across the CDM and SIDM simulations with different cross-sections by examining their residuals. Overall, the model performs similarly for CDM and the high-interaction SIDM scenario ($\sigma/m=10\,\text{cm}^{2}/\text{g}$), while showing lower coverage for the low-interaction SIDM case ($\sigma/m=1\,\text{cm}^{2}/\text{g}$). Despite some differences in distributions, the mean residuals for both $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}$ remain consistent within $1\sigma$ across simulations. These findings demonstrate the robustness of GraphNPE while highlighting the need for further testing with larger datasets to draw more definitive conclusions.
@@
In ongoing work, we aim to apply GraphNPE to observed dwarf galaxies in the Milky Way and the Local Group. We discuss key differences between galaxies in our mock FIRE-2 dataset and real observations in Section 7. Before applying GraphNPE to observational data, two key considerations need to be addressed: incorporating measurement uncertainties and masking contamination from background and binary stars. Both can be readily handled by integrating them into the training or inference process (see e.g. Wang et al. 2023).
In this work, we establish GraphNPE as a robust and efficient DF-based mapping method for inferring DM density profiles in dwarf galaxies, providing a promising avenue for constraining DM models. By leveraging machine learning, GraphNPE enables rapid, amortized inference, making it particularly well-suited for large-scale applications such as systematically exploring the mass modeling parameter space and investigating the effects of tidal disruption across mock simulations.
While this study applies GraphNPE to spherical equilibrium models, its framework is not inherently restricted to such cases, unlike traditional DF methods. GraphNPE can be extended to out-of-equilibrium systems (e.g. Ural et al., 2015; Widmark & Johnston, 2025) by expanding the training set to include galaxies undergoing tidal disruption, mergers, and other dynamical perturbations, which we will pursue in future work. This will allow GraphNPE to recover mass profiles in systems where equilibrium assumptions break down, further broadening its applicability to realistic astrophysical environments. More broadly, our work highlight the power of simulation-based inference and graph-based learning in astrophysics, paving the way for more adaptable and data-driven approaches to mass modeling in the era of increasingly precise kinematic surveys.
@@ Acknowledgments
We thank Jenna Samuel, Nondh Panithanpaisal, Nora Shipp, Stephanie O’Neil, Arpit Arora, Xiaowei Ou, Tjitske Starkenburg for helpful discussions.
TN, SM, and LN is supported by the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/). TN is also supported by a CIERA Postdoctoral Fellowship. LN is also supported by the Sloan Fellowship, the NSF CAREER award 2337864, NSF award 2307788. JIR would like to acknowledge support from STFC grants ST/Y002865/1 and ST/Y002857/1. CAFG was supported by NSF through grants AST-2108230 and AST-2307327; by NASA through grants 21-ATP21-0036 and 23-ATP23-0008; and by STScI through grant JWST-AR-03252.001-A. AW received support from NSF, via CAREER award AST-2045928 and grant AST-2107772. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics of U.S. Department of Energy under grant Contract Number DE-SC0012567.
@@ Appendix A Jeans dynamical modeling
$g(r)=\exp\left(2\int\frac{\beta(\tilde{r})}{\tilde{r}}\mathrm{d}\tilde{r}\right).$ (18)
@@ C.1 Comparison with Jeans-based methods
Figures 14 and 15 compare the inferred dark matter density profiles $\rho(r)$ for GraphNPE and Jeans modeling across the full set of selected FIRE galaxies in the CDM and SIDM simulations, respectively. Figures 16 and 17 show the velocity anisotropy profiles $\beta(r)$ for the same galaxies. Similarly, Figures 18 and 19 compare the inferred $\rho(r)$ profiles for GraphNPE and GravSphere in the CDM and SIDM simulations, respectively, while Figures 16 and 17 present the corresponding $\beta(r)$ profiles.
@@
C.2 Tidal disruption

### v2

@@ 6 Comparison between CDM and SIDM
Figure 13 shows the normalized distribution of fractional residuals for $V_{\mathrm{max}}$ (top) and $M_{\mathrm{200m}}^{\mathrm{peak}}$ (bottom), with CDM and SIDM simulations represented in purple and green, respectively. The mean and standard deviation of each distribution are indicated in the figure.
In the top panel, the mean residuals suggest that GraphNPE tends to overestimate $V_{\mathrm{max}}$ for CDM and SIDM1, consistent with the trends discussed in Section 5. This is further evidenced by the extended tails of the CDM and SIDM1 distributions. The model performs best on the SIDM10 samples, where the $V_{\mathrm{max}}$ residual distribution is more centered around zero. The residual distributions of $M_{\mathrm{200m}}^{\mathrm{peak}}$ in the bottom panel are more comparable between the simulations. The SIDM1 distribution peaks the highest at zero, but also exhibits some outliers toward positive residuals, contributing to a larger overall spread.
Table 6 presents the 68% and 95% coverage fractions for both $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$ across the CDM and SIDM samples. The reported means and associated errors are derived from 1000 bootstrap resamples. The model shows comparable performance on the CDM and SIDM10 simulations. In contrast, SIDM1 exhibits lower overall coverage, likely driven by the long tail of high residuals shown in Figure 6.
@@ 7 Discussion
Figure 15: The pericentric distance $r_{\mathrm{peri}}$ versus current distance $r_{\mathrm{curr}}$ for Milky Way dwarfs and our FIRE-2 mock dataset (gray contours). The contours are shown at every 10% intervals. Ultra-faint dwarfs $(M_{V}>-7.7)$ are denoted as squares. For clarity, dwarf galaxies within and outside of the 90% contour are denoted as filled and unfilled data points.
@@
Despite differences between the FIRE-2 mock samples and observations, it is important to note that the majority of the observed data samples remain well within the prior distribution. From Table 2, the Plummer scale radius can range from $r_{\mathrm{\star}}\in[20,20000]\,\mathrm{pc}$, [^13] For a Plummer profile, the projected half-stellar mass radius is equal to the scale radius, i.e., $R_{\star,1/2}=r_{\mathrm{\star}}$. while the velocity dispersion can range from $\sigma_{\mathrm{3-D}}\in[0.1,50]\,\mathrm{km/s}$ (or $\sigma_{\mathrm{los}}\in[0.06,28.9]\,\mathrm{km/s}$, assuming isotropy). The three galaxies with the smaller half-stellar mass radii are Triangulum II, Segue 1, and Willman 1 with values of $20.56^{+2.74}_{-2.61}\,\mathrm{pc}$, $23.90^{+3.74}_{-3.40}\,\mathrm{pc}$, and $27.43^{+5.89}_{-5.35}\,\mathrm{pc}$, respectively (see Pace 2024 and references therein). While these technically fall within the prior range, Triangulum II is very close to the lower boundary. This proximity may artificially truncate the posterior distribution, preventing a proper exploration of smaller values and potentially biasing parameter inferences. To better accommodate such observations, we plan to extend the prior limits in future analyses.
@@
Figure 15 shows the pericentric distance $r_{\mathrm{peri}}$ versus the current distance $r_{\mathrm{curr}}$ of the Milky Way dwarfs (circles) and the FIRE-2 simulations (gray contours, shown at every 10% interval. For clarity, we represent dwarfs within and outside of the 90% contour with filled and unfilled data points, respectively. Similarly, ultra-faint dwarfs $(M_{\mathrm{V}}>-7.7)$ are marked as squares and not expected to be resolved by FIRE-2. The vertical and horizontal dashed lines correspond to $0.2$ and $0.5$ times $R_{\mathrm{200m}}$ of the Milky Way, assumed to be $365\pm 76\,\mathrm{kpc}$ (Deason et al., 2020). Orbital properties are taken from Pace et al. (2022), which uses mixture models to determine the systemic proper motions of the Milky Way satellites and subsequently fits their orbits, accounting for both the Milky Way and LMC potentials. About half of the Milky Way satellites are within the 90% contour of the simulations, which typically occupy the high $r_{\mathrm{peri}}$ and $r_{\mathrm{curr}}$ regions. This difference highlights a numerical limitation: simulated satellites too close to their hosts are often disrupted and become much more challenging to track with rockstar (Behroozi et al., 2013; Mansfield et al., 2024).
@@ 7.2 Comparison with distribution function fitting
First, as with other NPE methods, GraphNPE is amortized, i.e. once trained, the model can be applied to any observation without re-training. In contrast, likelihood-based methods, including DF methods, must be rerun for each new dataset. This significantly limits the parameter space that can be explored with these models, particularly when extending to non-spherical cases. For systems with 1000 tracers, the Agama DF method in Read et al. (2021) takes approximately $\mathcal{O}(10^{3})$ minutes per run, comparable to the total training time of GraphNPE. However, once trained, GraphNPE performs inference in milliseconds, regardless of the tracer counts, making it far more efficient for large-scale applications. While computational cost may not be a major concern for real observations, given that the current number of observed dwarf galaxies does not yet pose a computational barrier (Pace, 2024), it becomes critical when testing and exploring the modeling parameter space on mock simulations. For instance, the detailed evaluations of mass modeling and tidal effects in Section 5 would be impractical with DF methods on the full FIRE-2 mock dataset, due to their high computational cost.
@@ 8 Conclusion
Dwarf galaxies and their mass density profiles hold the key to understanding structure formation and the particle nature of DM. Accurate mass modeling of these galaxies is essential for constraining DM properties and distinguishing between different theoretical scenarios, such as CDM and SIDM. As new spectroscopic surveys such as MUSE-Faint (Zoutendijk et al., 2020, 2021b, 2021a; Júlio et al., 2023) and DELVE (DECam Local Volume Exploration survey; Drlica-Wagner et al. 2022; Tan et al. 2025) continue to expand kinematic measurements of dwarf galaxy populations, robust modeling techniques that maximize the information extracted from stellar kinematics become increasingly valuable.
@@
  • When we compare GraphNPE to the GravSphere higher-order Jeans model, we find that the performance is overall similar, with discrepancies primarily arising from differences in modeling choices and prior distributions. Notably, the differences in prior distributions results in GravSphere (GraphNPE) favoring steeper (shallower) inner and outer density profiles.
@@
Lastly, we briefly compare the performance of GraphNPE across the CDM and SIDM simulations with different cross-sections by examining their residuals. Overall, the model performs similarly for CDM and the high-interaction SIDM scenario ($\sigma/m=10\,\text{cm}^{2}/\text{g}$), while showing lower coverage for the low-interaction SIDM case ($\sigma/m=1\,\text{cm}^{2}/\text{g}$). Despite some differences in distributions, the mean residuals for both $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$ remain consistent within $1\sigma$ across simulations. These findings demonstrate the robustness of GraphNPE while highlighting the need for further testing with larger datasets to draw more definitive conclusions.
@@
In ongoing work, we aim to apply GraphNPE to observed dwarf galaxies in the Milky Way and the Local Group. We discuss key differences between galaxies in our mock FIRE-2 dataset and real observations in Section 7. Before applying GraphNPE to observational data, several considerations need to be addressed: incorporating measurement uncertainties, accounting for selection biases, and masking contamination from foreground and binary stars. The first two can be readily handled by integrating them into the training or inference process (see e.g. Wang et al. 2023), while contamination requires careful data processing and member star selection.
An important direction for future work is extending GraphNPE to ultra-faint dwarfs, which represent some of the most DM-dominated systems in the universe. Although here we focus on classical dwarfs due to the FIRE-2 resolution limit, there is a compelling case that GraphNPE should perform well on the ultra-faints given their extremely low stellar tracer counts, a regime where GraphNPE has shown advantages over traditional methods. However, to properly validate GraphNPE for ultra-faint applications, we require higher-resolution cosmological simulations to stress-test and better quantify its performance on systems with low velocity dispersions and sparse kinematic data. As next-generation surveys continue to discover ultra-faint systems, establishing the reliability of mass modeling approaches in these data-limited regimes will be crucial for maximizing their scientific potential.
Furthermore, while this study applies GraphNPE to spherical equilibrium models, its framework is not inherently restricted to such cases, unlike traditional DF methods. GraphNPE can be extended to out-of-equilibrium systems (e.g. Ural et al., 2015; Widmark & Johnston, 2025) by expanding the training set to include galaxies undergoing tidal disruption, mergers, and other dynamical perturbations, which we will pursue in future work. This will allow GraphNPE to recover mass profiles in systems where equilibrium assumptions break down, further broadening its applicability to realistic astrophysical environments.
To conclude, in this work, we establish GraphNPE as a robust and efficient DF-based mapping method for inferring DM density profiles in dwarf galaxies, providing a promising avenue for constraining DM models. By leveraging machine learning, GraphNPE enables rapid, amortized inference, making it particularly well-suited for large-scale applications such as systematically exploring the mass modeling parameter space and investigating the effects of tidal disruption across mock simulations. More broadly, our work highlights the power of SBI and graph-based learning in astrophysics, paving the way for more adaptable and data-driven approaches to mass modeling in the era of increasingly precise kinematic surveys.
@@ Acknowledgments
We thank Jenna Samuel, Nondh Panithanpaisal, Nora Shipp, Stephanie O’Neil, Arpit Arora, Xiaowei Ou for helpful discussions.
TN, SM, and LN is supported by the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/). TN is also supported by a CIERA Postdoctoral Fellowship. LN is also supported by the Sloan Fellowship, the NSF CAREER award 2337864, NSF award 2307788. JIR would like to acknowledge support from STFC grants ST/Y002865/1 and ST/Y002857/1. CAFG was supported by NSF through grants AST-2108230 and AST-2307327; by NASA through grants 21-ATP21-0036 and 23-ATP23-0008; and by STScI through grant JWST-AR-03252.001-A. AW received support from NSF, via CAREER award AST-2045928 and grant AST-2107772. TS gratefully acknowledges the support of the NSF-Simons AI-Institute for the Sky (SkAI) via grants NSF AST-2421845 and Simons Foundation MPS-AI-00010513 and support by NASA grants 22-ROMAN22-0055 and 22-ROMAN22-0013. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics of U.S. Department of Energy under grant Contract Number DE-SC0012567.
@@ Appendix A Jeans dynamical modeling
[…]
@@ C.1 Comparison with Jeans-based methods
Figures 16 and 17 compare the inferred DM density profiles $\rho(r)$ for GraphNPE and Jeans modeling across the full set of selected FIRE galaxies in the CDM and SIDM simulations, respectively. Figures 18 and 19 show the velocity anisotropy profiles $\beta(r)$ for the same galaxies. Similarly, Figures 20 and 21 compare the inferred $\rho(r)$ profiles for GraphNPE and GravSphere in the CDM and SIDM simulations, respectively, while Figures 18 and 19 present the corresponding $\beta(r)$ profiles.
@@
C.2 Performance as a function of tracer counts
We examine the performance of GraphNPE as a function of tracer counts, $N_{\star}$. Figure 24 presents the absolute residuals of $V_{\mathrm{max}}$ (left panel) and $M_{\mathrm{200m}}^{\mathrm{peak}}$ (right panel) across four $N_{\star}$ bins with edges at $[20,50,100,1000,5000]$. We divide the samples into FIRE-2 CDM and SIDM galaxies and show the result separately as purple and green data points.
The absolute residuals exhibit a clear decline with increasing tracer count $N_{\star}$. This trend is similar to that observed with $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$ in Figures 11 and 12, which is expected given the strong correlations between $N_{\star}$ and the galaxy mass (e.g. the stellar-to-halo-mass relation). For $V_{\mathrm{max}}$, residuals are approximately $20\%$ at low tracer counts ($N_{\star}<100$), decreasing to a plateau of roughly $10\%$ for $N_{\star}>100$. For $V_{\mathrm{max}}$, residuals are approximately $20\%$ at low tracer counts ($N_{\star}<100$), decreasing to a plateau of roughly $10\%$ for $N_{\star}>100$. Notably, both $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$ exhibit considerably larger residual scatter in the low tracer count regime. Finally, we observe that GraphNPE exhibits marginally superior performance on SIDM galaxies at high tracer counts and on CDM galaxies at low tracer counts, though these differences are small compared to the residual scatter.
Figure 24: The absolute residuals of $V_{\mathrm{max}}$ (left) and $M_{\mathrm{200m}}^{\mathrm{peak}}$ (right) as a function of the tracer counts. FIRE-2 CDM galaxies are shown as purple data points, while SIDM galaxies are shown as green data points. The vertical error bars denote the standard deviation of the absolute residuals in each bin. The horizontal error bars denote the bin width.
C.3 Tidal disruption